# Autodeleveraging: Impossibilities and Optimization

### v1

@@ I.2 Connection to Classical Impossibility Results
    Probabilistic CAP Theorem: Blockchains circumvent the strict CAP theorem by weakening consistency to probabilistic finality (e.g., Nakamoto consensus) or availability to liveness under synchronous periods [PassShi2017, Shi2020, ShiConsensusBook]. This parallels the (S) vs. (R) trade-off: exchanges effectively accept probabilistic solvency (via insurance funds) to maintain liveness (continuous trading/revenue).
@@
    Cryptographic Commitments for Credibility: The credibility trilemma of Akbarpour and Li motivates cryptographic mechanisms that make optimal auctions simultaneously credible and strategy-proof by enforcing operator commitments [AkbarpourLi2020, FerreiraWeinberg2020, EssaidiFerreiraWeinberg2022, ChitraFerreiraKulkarni2024]. For ADL, this suggests that verifiable execution (e.g. via zero-knowledge proofs or on-chain logic) could allow an exchange to commit to a dynamic policy that balances the trilemma better than any opaque static policy could, by removing the operator’s incentive to deviate during crises.

### v3

@@ I.3 Diagnostic decomposition (not the §9 overshoot metric)
    Probabilistic CAP Theorem: Blockchains circumvent the strict CAP theorem by weakening consistency to probabilistic finality (e.g., Nakamoto consensus) or availability to liveness under synchronous periods [92, 104, 105]. This parallels the (S) vs. (R) trade-off: exchanges effectively accept probabilistic solvency (via insurance funds) to maintain liveness (continuous trading/revenue).
@@
    Cryptographic Commitments for Credibility: The credibility trilemma of Akbarpour and Li motivates cryptographic mechanisms that make optimal auctions simultaneously credible and strategy-proof by enforcing operator commitments [7, 54, 53, 34]. For ADL, this suggests that verifiable execution (e.g. via zero-knowledge proofs or on-chain logic) could allow an exchange to commit to a dynamic policy that balances the trilemma better than any opaque static policy could, by removing the operator’s incentive to deviate during crises.